OpenStudy (anonymous):

Whats the slope of (1,1) (6,3)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

by definition slope m = (y-y1)/(x-x1) so, in your case: m=(3-1)/(6-1)=2/5

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) m=(3-1)/(6-1) m=2/5

OpenStudy (anonymous):

Thank you so much :D

OpenStudy (anonymous):

you are welcome. :)

OpenStudy (anonymous):

Now if only I had a cookie. lol

OpenStudy (anonymous):

lol, why not?

OpenStudy (anonymous):

thank you for that mrs. george.

OpenStudy (anonymous):

Your welcome ;)

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

lol